What is the firm's profit? (Round to two decimal points)

Assume a competitive firm faces a market price of $120, a cost curve of C20q 500, and a marginal cost of MCqt 20 What is the firms profit maximizing output level? 10 units. (round your answer to two decimal places) What is the firms profit maximizing price? $ 120.(round to the nearest penny) What is the firms profit? S(round to the nearest penny)
Answer #1

Since the firm is the competitive firm price will be same for all levels of production since these firms sell homogeneous products. Here Price = Average revenue = Marginal revenue.

Profit maximizing condition is where marginal revenue = marginal cost. Here marginal revenue = price = $120.

Marginal revenue = Marginal cost

120 = q^2 + 20

q^2 = 100

q = 10 (profit maximizing level of output)

Profit maximizing price level is $120.

Profit = Total revenue - total cost

Total revenue = Price*Quantity

= 120*10 = 1200

Total cost = 1/3 q3 + 20q + 500

TC = 1/3 (10)3 + 20(10) + 500

TC = 333.33 + 200 + 500

= 1033.33

Profit = Total revenue - total cost

= 1200 - 1033.33

= $166.67